In a goodness of fit test of n=100 trials, if the no. of different categories or cells k = 3 and H0: p1=0.2; p2=0.3; p3 =0.5 then expected frequency for each cell is equal to 20, 30 and 50 respectively.

T OR F?
Answer #1

Solution:

Given:

n = number of trials = 100

k = number of cells = 3

H0: p1=0.2; p2=0.3; p3 =0.5

Expected frequencies are calculated as:

Ei = n * pi

i = 1 ,2,3

E1 = n * p1 = 100 * 0.2 = 20

E2 = n * p2 = 100 * 0.3 = 30

E3 = n * p3 = 100 * 0.5 = 50

thus given statement : expected frequency for each cell is equal to 20, 30 and 50 respectively is TRUE.

Thus answer is: T